PLUS Help IC College Physics (PHYC 2053 & 205 Chapter 02, Problem 06 One afternoon, a couple walkst the distance they travel? (b) What is the magnitude of the couples displacement? (c) What is the direction (relative to due east) of the couples displacement? of the way around a circular lake, the radius of which is 1.74 ikm. They start at the west side of the lake and head due south to begin with. (a) What is MacBook Air
Answer #1

answer) a) the distance they travel=3/4*C

C=circumference of a circle=2pir

so d=3/4*2pir=3/2*pi*1.74km=8.20 km

so answer is 8.20 km

b) from the triangle we can see that

\Deltax=xf-xi

\Deltax=\sqrt{}r2+r2=\sqrt{}2r2=(\sqrt{}2)* r=(\sqrt{}2)*1.74=2.46km

so the answer is 2.46 km or 2.5 km

c) from the triangle we can see

tan\Theta=r/r

tan\Theta=1

\Theta=tan-1(1)=45o

so the answer is 45o north of east.